### v2

# OORD: The Oxford Offroad Radar Dataset

Abstract
There is a growing academic interest as well as commercial exploitation of millimetre-wave scanning radar for autonomous vehicle localisation and scene understanding. Although several datasets to support this research area have been released, they are primarily focused on urban or semi-urban environments [1, 2, 3, 4]. Nevertheless, rugged offroad deployments are important application areas which also present unique challenges and opportunities for this sensor technology. Therefore, the Oxford Offroad Radar Dataset (OORD) presents data collected in the rugged Scottish highlands in extreme weather. The radar data we offer to the community are accompanied by GPS/INS reference – to further stimulate research in radar place recognition. In total we release over of radar scans as well as GPS and IMU readings by driving a diverse set of four routes over forays, totalling approximately of rugged driving. This is an area increasingly explored in literature, and we therefore present and release examples of recent open-sourced radar place recognition systems and their performance on our dataset. This includes a learned neural network, the weights of which we also release. The data and tools are made freely available to the community at oxford-robotics-institute.github.io/oord-dataset
Index Terms:
Radar, Place Recognition, Localisation, Odometry, Positioning, Robotics, Autonomous Vehicles
I Introduction
Compared to cameras and LiDARs, radar systems offer distinct advantages, including extended-range perception and resilience to lightning and various weather conditions. In more detail, scanning radar of the type we use in this work does not give 3D information, and is suspect to complex measurement artefacts (speckle noise, ghost reflections, etc), but is extremely robust to atmospheric effects and low-light. Cameras are information-rich and highly discriminative, but are highly susceptible to appearance variation, e.g. low-light. LiDARs give detailed scans, often in 3D, and emit their own light so are suitable for low-light operation, but are often limited in range and are nevertheless susceptible to particulate effects, e.g. snow or dust. Full reviews of millimetre-wave radar applications in autonomous vehicle tasks are available in e.g. [5, 6]. In brief, scanning radar-based motion estimation [7, 8, 9, 10, 11, 12, 13, 14], pose estimation [15, 16, 17], place recognition [18, 19, 15, 20, 21, 22, 23, 24, 25, 26, 27], and full simultaneous localisation and mapping (SLAM) [28, 29] are increasingly popular in the literature. This dataset is designed to support that work, and to prompt an extension into naturalistic, rugged environments, which have not yet been explored.
Natural environments are important because off-road autonomous vehicles have many essential applications across industries operating on rugged terrain. This technology is essential for tasks such as agricultural operations, search and rescue missions, and exploration in remote areas, where human intervention may be impractical or hazardous. Also, autonomous hauling or resource extraction is important in mining, autonomous tractors for planting and harvesting can optimise crop yields, tree planting and logging can reduce environmental impact and risk to workers, autonomous grading and excavation can enhance efficiency and safety in construction, search and rescue missions can leverage autonomous drones to locate missing persons in hazardous locations, and autonomous vehicles can be of great use in environmental monitoring and disaster response.
In deploying autonomous machines to these sites, place recognition is vital for navigation and localisation. For this, more robust sensors are crucial to enhancing place recognition accuracy and reliability. Radar in particular can handle challenging conditions like low-light environments.
We therefore present an offroad radar dataset with a focus on relocalisation, with Fig. 1 showing an example location from our collection site. Our contributions are as follows:
1.
Offroad scanning radar data We release the first radar dataset focusing on off-road, difficult terrain and naturalistic environments,
2.
Poor weather and lighthing conditions including collection after thick snowfall and in total darkness (being in the wilderness),
3.
Place recognition ground truth Our raw data is referenced against good GPS as a ground truth for the place recognition task,
4.
Open-sourced radar place recognition implementations For the first time, we release open-source a set of trained weights for a deep neural network which solves the radar place recognition,
5.
Software development kit We release software tools for making the use of this dataset in some pre-existing, non-learned open-source radar place recognition implementations easy, as well as pipelines to accelerate new learning solutions.
In Sec. III we describe and justify the choice of collection sites. Sec. IV describes the data collection platform, sensor format including radar as well as the GPS/INS synchronised with this main sensor, and the download data structure for how to interact with the dataset. Sec. V presents several useful software tools which make it easy to apply this data, e.g. with common performance metrics and data formats. Finally, Sec. VI shows examples based off of the tools in Sec. V for several publicly available place recognition algorithms as tested against this dataset.
Drone image of route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
DSLR image from route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
DSLR image from route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
DSLR image from route.
GPS trace (red).
Ground truth matrix.
Example scan #1.
Example scan #2.
Example scan #3.
Example scan #4.
| Dataset | Variable terrain | Poor weather | Scene diversity | Repeat traversals |
| Radar RobotCar | ✗ | ✗ | ✗ | ✓ |
| Marulan | ✓ | ✓ | ✓ | ✓ |
| MulRan | ✗ | ✗ | ✓ | ✓ |
| RADIATE | ✗ | ✓ | ✓ | ✗ |
| Boreas | ✗ | ✓ | ✓ | ✓ |
| OSDaR23 | ✗ | ✗ | ✓ | ✗ |
| OORD (Ours) | ✓ | ✓ | ✓ | ✓ |
II Related Work
II-A Radar datasets
Datasets featuring a similar class of millimetre-wave scanning radar include Marulan [30] by Peynot et al, which is furthermore also collected in natural environments, as is ours. This data, however, only has range of up to . Several recent radar datasets use the same radar class – from the same manufacturer111Navtech Radar: navtechradar.com – as used in our work, including the Oxford Radar RobotCar Dataset [31] by Barnes et al, MulRan [18] by Kim et al, RADIATE [32] by Sheeny et al, Boreas [33] by Burnett et al, and OSDaR23 [34] by Tagiew et al. The Oxford Radar RobotCar Dataset [31] is purely urban and features many repeat traversals, useful in investigating place recognition. The ground truth provided in [31], however, is prepared for radar odometry rather than place recognition. MulRan [18] presents more diversity in scenery, having been collected in several distinct parts of Daejeon, South Korea. We similarly attend to generalisation-sensitive training/testing data and collect at many sites, albeit offroad and more widespread across a large area of countryside/wilderness rather than a city. Our repeat traversals are also more numerous than for MulRan. RADIATE [32], presenting vehicle detection and tracking ground truth, is similarly lacking in overlapping traversals and not geared towards place recognition, although a diversity in weather makes this dataset interesting for exploring this sensor’s robustness. Closest to our work is Boreas by Burnett et al [33], who collect multi-season data including a Navtech CIR304-H scanning radar. Our work is in part inspired by this, with ours featuring highly non-planar offroad driving as opposed to purely urban.
Tab. I summarises the characteristics of our dataset as compared to these.
II-B Radar place recognition
Owing to its inherent resilience in sensing capabilities, radar has attracted considerable attention for its adeptness in the face of adverse weather, demonstrating a capacity to navigate and operate effectively in inclement conditions. This is because radar systems typically use electromagnetic waves in the radio or microwave frequency range. Unlike visible light, these waves can penetrate through various weather elements like rain, snow, fog, and clouds without significant distortion or attenuation. Radar place recognition has thus been explored in [19, 15, 35, 20, 22, 23, 36, 37, 38, 27].
This has included simultaneous localisation and mapping [36, 37] with loop closure detection mechanisms, where for example in [36] loop candidates are rejected by principal components analysis to focus on distinct structural layouts.
There are also several methods for learning to vectorise radar scans with neural networks [19, 15, 23, 20, 22, 27], including supervised [19, 15, 20, 20] and self-supervised approaches [27, 23] and learned uncertainty [27].
Finally, some approaches are non-learned [18, 35, 38, 39], instead relying largely on transformation-vinvariant integral transforms (e.g. Fourier, Fourier-Mellin, Radon) as in [38, 39], pooling operations such as max/min/mean to achieve rotation invariance and to reduce the representation dimensionality [18, 39], and circular cross-correlation for an alignment/similarity score [18, 38].
This place recognition capability can be crucial for simultaneous localisation and mapping (SLAM), where reinitialisation is essential for correcting errors and maintaining accuracy over time. SLAM is a newly developed area for scanning radar in autonomous vehicle operations, as per [28, 29, 37], and so robust reinitialisation through radar place recognition is an important research topic. This area is therefore of increasing interest, but far from mature, and this dataset should therefore contribute to establishing useful benchmarks for the community.
III Off-road routes & collection sites
Figs. 3 and 5 show a detailed overview of the datasets we provide, at various areas of the Ardverikie Estate. Data has been gathered in the region encompassing Loch Laggan and Lochan na h-Earba. The position of these lochs can be particularly well understood by referring to the caption of Fig. 5(b).
Loch Laggan, a freshwater loch located about west of Dalwhinnie in the Scottish Highlands, extends in a nearly northeast to southwest direction for approximately . The name Lochan na h-Earba actually refers to two lochs positioned south of Loch Laggan. These lochs are situated in a slender glen that runs from southwest to northeast, running roughly parallel to Loch Laggan. The Binnein Shuas range of hills separates them from Loch Laggan.
Each of Figs. 5(b), 3(a), 5(a) and 3(b) is arranged as follows: Top Left: Photograph of each dataset collection location (handheld camera, not on-board imagery). Top Middle: GPS trace of route driven. Top Right: Ground truth match matrix for a pair of trajectories at that site. Bottom: Example radar scans from various points along the route (Cartesian).
We release data from four areas of the Ardverikie Estate. These feature distinct landscape (therefore typical radar returns) as well as driving conditions. We refer to the sites as Two Lochs, Bellmouth, Maree, and Hydro. Each of these is discussed below. Tab. II summarises the collection sites and routes. Fig. 6 compares the elevation along each route.
| Route | Distance | # Radar | # Camera | # GPS | Size |
| Bellmouth | |||||
| Hydro | |||||
| Maree | |||||
| Two Lochs |
III-A Route 1: Bellmouth
This is a route. The route is completely landlocked, heading away from and then back towards Loch Laggan, see Fig. 3(a) (Top Middle. The ascent and descent are reflected by a somewhat symmetrical elevation change in Fig. 6 – climbing for the first half of the route, and descending for the second half. The vehicle is always on a gravel track. Figs. 7(b), 7(a), 7(d) and 7(c) show images from the vehicle’s perspective from four outings along this route. The first two these outings is prior to any snow falling in the Ardverikie Estate area (Figs. 7(a) and 7(b)). The next two of these outings show significant snow dump (Figs. 7(c) and 7(d)). The outings are in the same direction, as per the ground truth matrix in Fig. 5(b) (Top Right), but within each outing itself we double back (c.f. the off diagonal matches in three parts of this matrix).
III-B Route 2: Hydro
This is a route. As shown in Fig. 6 this route reaches the highest elevation of all or our collection sites – ascending a gravel track to maximum elevation and then descending along the same route. For parts of the route, we travel along a mountain river, as per Fig. 3(b) (Top Middle). The snow coverage is significant (Figs. 7(e), 7(f) and 7(g)), but not as thick as for Maree below.
III-C Route 3: Maree
This is a route. The route is mountainous, in the elevated area to the south of Loch Laggan, see Fig. 5(a) (Top Middle). There are two ascents and descents as can be seen in Fig. 6 – with a descent halfway through the journey, and another at the end of the journey. The vehicle is primarily on a gravel track. Figs. 7(h) and 7(i) show images from the vehicle’s perspective from two outings along this route. Both of these feature heavy snow, with a completely covered track. In Fig. 7(i) the vehicle is driving in total darkness (with no artificial illumination beyond the vehicle’s spotlights as this is a very remote and isolated natural collection site). The outings are in the same direction, as per the ground truth matrix in Fig. 5(a) (Top Right), but within each outing itself we double back on parts of the route and return to the starting point along the same route we took to begin the route (c.f. the off diagonal matches in three parts of this matrix).
III-D Route 4: Two Lochs
This is a route across the entire estate. The route is alongside two lochs (Loch Laggan and Lochan na h-Earba), see Fig. 5(b) (Top Middle, where the lochs are on top of and the bottom of the image, respectively). Referring to the route in an anti-clockwise sense, while going alongside Loch Laggan the vehicle is on tarmac – it is otherwise on unpaved road for the rest of the route. Figs. 7(k) and 7(j) show images from the vehicle’s perspective of our two outings – these are completed after snow has fallen (in previous days), so melted somewhat, but while it is still very visible on the ground etc. The outings are in the opposite direction as shown by the off-diagonal matches in the ground truth matrix in Fig. 5(b) (Top Right). Note that the red elevation profile in Fig. 6 corresponds to the foray shown in Fig. 5(b) (Top Middle), and is very flat at approximately as the vehicle traverses alongside Lochan na h-Earba.
III-E Summary: Datasets provided
In summary, the datasets provided are listed below, including the date-string for the collection time and the site that they were collected at, with colours corresponding to Fig. 6.
[…]
(Bellmouth #1)[DB1]: 2021-11-25-12-01-20
[…]
(Bellmouth #2)[DB2]: 2021-11-25-12-31-19
[…]
(Bellmouth #3)[DB3]: 2021-11-26-15-35-34
[…]
(Bellmouth #4)[DB4]: 2021-11-26-16-12-01
[…]
(Hydro #1)[DH1]: 2021-11-27-14-37-20
[…]
(Hydro #2)[DH2]: 2021-11-27-15-24-02
[…]
(Hydro #3)[DH3]: 2021-11-27-16-03-26
[…]
(Maree #1)[DM1]: 2021-11-28-15-54-55
[…]
(Maree #2)[DM2]: 2021-11-28-16-43-37
[…]
(Two Lochs #1)[DA1]: 2021-11-29-11-40-37
[…]
(Two Lochs #2)[DA2]: 2021-11-29-12-19-16
These are shown in Fig. 7. We note that, beyond the weather and illumination effects discussed for each route above, some forays have lens adherents on the on-board camera (e.g. Figs. 7(c) and 7(d) in particular), motivating the use of radar in this sort of domain and for the localisation task.
IV Sensors & dataset format
At oxford-robotics-institute.github.io/oord-dataset we provide download time-stamped synchronised radar and GPS/INS.
IV-A Radar
This sensor is the same as that used for the Oxford Radar RobotCar Dataset [31], i.e. a Navtech CTS350-X Millimetre-Wave frequency-modulated continuous wave radar, , measurements per rotation, range, range resolution, beamwidth. The radar was mounted at the centre of the vehicle aligned to the vehicle axes.
Radar scans are stored as lossless-compressed PNG files in polar form with each row representing the sensor reading at each azimuth and each column representing the raw power return at a particular range.
IV-B Camera
Our vehicle was also equipped with a FLIR Blackfly USB3 Vision camera222flir.com/products/blackfly-s-usb3 which captures images at resolution over a sensing area and with of dynamic range.
Some example images are shown in Fig. 7, and we present experimental results comparing vision to radar in Sec. VII. Note that in Tab. II the camera framerate has been downsampled to more approximately match the radar frame rate (and mitigate download sizes).
This will be useful for sensor fusion research such as [40], where our radar and camera are timestamp-synchronised.
IV-C GPS/INS
We use a Microstrain 3DM-RQ1-45 GPS/INS. This sensor has direct satellite and inertial measurements, and computes position, velocity, and attitude. To provide accurate readings, it uses a triaxial accelerometer, gyroscope, magnetometer, and temperature sensors, as well as a pressure altimeter. Furthermore, dual on-board processors run a Extended Kalman Filter (EKF).
IV-D Calibration
Extrinsic calibration between the sensors is done by reference to a detailed CAD model of the vehicle and sensor suite. These are included in the SDK (Sec. V) as an extrinsics.yaml file specifying the relative six degree-of-freedom pose between each of the GPS/INS (Sec. IV-C)
which is to say that the GPS unit is approximately behind the radar with axes aligned to the radar. Then, for the camera (Sec. IV-B) to radar
which is to say that the camera is further behind and to the right of the radar with axes offset by two rotations. Finally, the radar (Sec. IV-A, as the main sensor focused on in this work) is at , i.e. is used as the datum/origin of the frame of reference.
V Software & experimental tools
At github.com/mttgdd/oord-dataset and huggingface.co/mttgdd/oord-models we provide a software tools to support the use of this dataset.
V-A Downloads & data loaders
Similarly to [31], we provide data loaders to work with the radar data format, which as per Sec. IV-A are released in raw polar format, but can be easily transformed into Cartesian format with our SDK.
For working with driven sequences, we provide a basic iterable Dataset class which returns each radar scan paired with its closest (in a timestamp sense) GPS/INS reading (Sec. IV-C).
We also provide a batch download script to rapidly access the data.
V-B Radar place recognition & evaluation
In place recognition it is common to represent the input image or scan by some descriptor vector. With map descriptors at places indexed by and of dimension and a query descriptor , place recognition can be stated as
where we decide that we are in place given that the descriptor for that place is closest to our query descriptor (i.e. “place recognition”). Alternatively, instead of Eq. 3, we may search for a set of nearest neighbours.
We evaluate place recognition performance by comparing “distance matrices” in the ground truth space given by GPS, as shown by example in Figs. 3 and 5 (Top Right) to distance matrices composed of e.g. pairwise distances between vectors representing radar scans (see Sec. VI) or some other similarity measurement. These matrices have heights given by the number of radar scans in a “query” foray and widths similarly for a “reference” foray.
Specifically, with respect to Eq. 3 above we have black and white distance matrices as inin Figs. 3 and 5 (Top Right) given by
where is the UTM northing/easting GPS coordinates corresponding to the query/localisation trajectory scan and reference/map trajectory scan , with and with and as the number of frames in these trajectories, respectively. Similarly, we have
with and embeddings of the query/reference radar scans at those same locations . After reducing to its binarised form with in each row indicating the nearest neighbour (or, generally, a set of s indicate some nearest neighbours) as per Eq. 3, we expect for good place recognition that (with a numerical accuracy defined as Recall@N below).
Highlighted regions are such that those radar scan pairs have locations given by GPS which are within (this threshold being configurable in our SDK). In these, interesting features to note are the off-diagonal revisits, which represent traversals in the opposite direction. This is especially important in the radar modality as the sensor has a field-of-view.
Then, Recall@N refers to the percentage of query scans which result in at least one successful localisation (i.e. resulting in a returned location within of the query location) when accepting as candidates all of the nearest neighbours in the embedding space. This is a commonly used metric and threshold, as in visual place recognition [41, 42].
Consider the black and white matrices shown in Figs. 5(b), 3(a), 5(a) and 3(b) (Top Right). Each of these has, as per Eq. 4 and Eq. 5, a number of rows equal to the number of scans (with some examples in the bottom row of Figs. 3 and 5) collected along a query route (shown by the red GPS traces in Figs. 3 and 5) and number of columns equal to the number of scans collected along a reference route – i.e. the reference route acts as a map and the query route is localised against it. Each cell in these matrices are thus a pair of scans from a pair of locations. The black background means that the distance between those locations is and white entries indicate that the pair of locations is . Note that in Figs. 3 and 5 the true positives are “thickened” for visualisation purposes, to , but we experiment in Sec. VII with .
For this we use a “difference matrix” in Fig. 8 with the embedding distances (Euclidean) between live and map features. Not that the “grids” apparent in Figs. 8(a) and 8(b) are artefacts resulting from embeddings representing scans being very similar to many such embeddings in the map. These are evident as “blocks” of blue in the difference matrix. Within this block, all embeddings are very similar (where yellow indicates embeddings being more dissimilar). This phenomena is therefore known as “perceptual ambiguity” or “aliasing”. This indicates poor representational capability of the method, where we want only scans from the same location to be similar – see Figs. 8(c) and 8(d), in which only the diagonal and lines perpendicular to the diginal are blue, with no such gridlike structure. Localisation is successful if the nearest neighbour for a query embedding (rows) is a reference embedding (column) which is in truth (c.f. the ground truth matrices in Fig. 3(a) Top Right) is close to the query location in physical space.
V-C Radar place recognition with neural networks
As introduced in Sec. V-B, radar place recognition is concerned with learning discriminative and compressed representations of radar scans, to maximise similarity between likewise places and minimise similarity otherwise. However, for this and similar sensors – despite their inherent immunity to appearance change – the beam is not narrow, power returns are perturbed by noises (e.g. speckle) and artefacts (multi-path reflections, etc). However, radar data – despite being complex in this way – contains highly informative information. While it is difficult to manually/visually inspect radar data for this information, it has been found in [19, 35, 27, 23] that neural networks are able to learn to encode very informative feature representations that are germane to the task of place recognition. As will be discussed below, we evaluate a family of learned methods on our dataset. This includes a pretrained and frozen backbone feature extractor along with a pooling layer which is fine-tuned on radar data from this dataset, and we release these weights at huggingface.co/mttgdd/oord-models, with this being, to our knowledge, the first open-sourced learned radar place recognition system.
V-D Open-source radar place recognition methods
As below in Sec. VI, our SDK includes either (1) new implementations of or (2) imports, as submodules, of other open-source radar place recognition systems [18, 26, 39]. Experiments on novel systems can be quickly run over all pairs of trajectories discussed in Sec. III-E, and prepared in the format of Tab. III, by a common .yaml configuration format.
VI Examples
We demonstrate the utility of our dataset as a novel radar place recognition benchmark using the evaluation method described in Sec. V-B. We also show that the GPS/INS described in Sec. IV-C can be put to good use in supervising learned radar place recognition methods. Therefore, we demonstrate implementations of the following methods with our dataset , including four methods, LABEL:B1–LABEL:B4, which are either not fine-tuned at all (by virtue of the method itself) or are fine-tuned on this data as well as four methods which are totally pre-trained, LABEL:B5–LABEL:B9, i.e. not fine-tuned on this new dataset.
Non-learned radar-specific methods
[…]
(Baseline 1)[B1]: RingKey as a component of ScanContext [43] averages point cloud contents at a fixed distance around the vehicle (thus being rotation invariant). Note, for this we do not use the orientation refinement of ScanContext, as distances between row-reduced vectors and orientation scores are separate, i.e. these form a hierarchy of localisers. To be clear, this is different to LABEL:B2 below, where the maximum circular cross-correlation is the score used for similarity between scans (with no initial score used to reduce the candidate set).
[…]
(Baseline 2)[B2]: RaPlace [26] measures the similarity score between radar scans by using Radon-transformed sinogram images and cross-correlation in the frequency domain. This gives rigid transform invariance during place recognition, and supresses the effects of radar multipath and ring noises.
[…]
(Baseline 3)[B3]: Open-RadVLAD [39] uses only the polar representation. Also, for partial translation invariance and robustness to signal noise, it uses only a 1D Fourier Transform along radial returns. It also achieves rotational invariance and a very discriminative descriptor space by building a vector of locally aggregated descriptors (VLAD).
Fine-tuned radar-specific methods
[…]
(Baseline 4)[B4]: Here we train radar-specific neural network models – using ResNet18 [44] as a backbone feature extractor and NetVLAD [41] as a pooling layer (with clusters). Inputs are polar radar scans of resolution and embeddings are -dimensional. The radar returns are replicated times at the input in order to form a -channel image. We start with pretrained weights learned against ImageNet [45] and then, in an unsupervised learning step, initialise the NetVLAD cluster centres by k-means++ [46] over the deep features extracted from the reference trajectory radar frames.
Pre-trained methods
[…]
The pretrained networks are selected as they have been applied to visual place recognition in other work – e.g. MobileNet by Wu et al. [51], AlexNet by Sünderhauf et al. [52], GoogleNet by Yue et al. [53], VGG by Arandjelovic et al. [41], and ResNet by Yu et al. [54] – and it is therefore important to apply them to place recognition in the radar domain. Sec. VII presents a comprehensive comparison of these example baselines over all pairs of trajectories from our dataset. A comparison of the computational efficiency of these various baselines is outside of the scope of this work, but the interested reader is referred to [55] (Figure 2, page 2) for an in-depth comparison of inference speeds and memory requirements for the pretrained networks LABEL:B5–LABEL:B9 and to [39] for a comparison of the radar-specific methods, LABEL:B1–LABEL:B3, where Open-RadVLAD is shown to be more computationally efficient than RaPlace.
| LABEL:DB2-vs-LABEL:DB1 | LABEL:DH1-vs-LABEL:DH2 | LABEL:DH1-vs-LABEL:DH3 | LABEL:DM1-vs-LABEL:DM2 | LABEL:DA2-vs-LABEL:DA1 | |
| LABEL:B1 | 94.29 / 97.53 / 98.24 | 91.15 / 93.57 / 93.78 | 94.06 / 95.80 / 96.00 | 91.76 / 97.22 / 98.42 | 72.96 / 82.28 / 84.71 |
| LABEL:B2 | 98.45 / 98.73 / 98.94 | 92.65 / 93.89 / 93.94 | 95.34 / 95.85 / 95.85 | 94.28 / 99.15 / 99.40 | 85.82 / 87.69 / 88.25 |
| LABEL:B3 | 98.87 / 99.79 / 99.93 | 93.51 / 94.10 / 94.16 | 96.00 / 96.16 / 96.16 | 96.88 / 99.06 / 99.32 | 84.27 / 87.36 / 88.25 |
| LABEL:B4 | 97.67 / 98.52 / 98.87 | 93.78 / 94.26 / 94.37 | 95.70 / 99.59 / 99.95 | 97.14 / 99.06 / 99.44 | 5.41 / 10.98 / 16.23 |
| LABEL:B5 | 90.49 / 97.25 / 98.10 | 92.76 / 93.83 / 93.83 | 94.72 / 95.85 / 95.85 | 89.11 / 97.91 / 98.72 | 4.53 / 10.10 / 15.51 |
| LABEL:B6 | 90.70 / 97.04 / 98.45 | 91.96 / 93.62 / 93.83 | 93.14 / 95.49 / 95.75 | 87.92 / 96.67 / 98.25 | 17.55 / 40.07 / 50.44 |
| LABEL:B7 | 94.50 / 98.38 / 98.80 | 92.28 / 93.83 / 93.83 | 93.70 / 95.54 / 95.80 | 90.61 / 97.27 / 98.42 | 23.18 / 47.19 / 58.22 |
| LABEL:B8 | 74.91 / 92.39 / 95.28 | 82.79 / 91.31 / 92.60 | 83.91 / 92.78 / 94.62 | 65.16 / 86.93 / 92.10 | 14.74 / 35.87 / 47.08 |
| LABEL:B9 | 72.73 / 90.49 / 94.71 | 81.29 / 90.40 / 91.96 | 84.07 / 93.14 / 94.31 | 69.43 / 87.15 / 91.97 | 9.71 / 21.80 / 31.29 |
VII Results
VII-A Baseline comparison
As can be seen in Tab. III, considering that all of each of the routes is at minimum in length (seeTab. II), all radar-specific methods perform very well, with in excess of for even Recall@1 i.e. when relying on only a single nearest neighbour in embedding space look up. Tab. III also shows results when using and nearest neighbours. Additionally, Fig. 9 shows results when using all of to nearest neighbours, as “recall curves” rather than tabulated.
Overall the best performing methods are LABEL:B3 Open-RadVLAD [39], LABEL:B4 ResNet18-NetVLAD, and LABEL:B2 RaPlace [26].
Fig. 9 shows the Recall@N curves corresponding to Tab. III, averaged over all pairs of trajectories. From this we can see that LABEL:B3 Open-RadVLAD and LABEL:B4 ResNet18-NetVLAD perform best with LABEL:B3 Open-RadVLAD slightly better at lower recalled candidates (and thus perhaps being better suited to efficient deployment) but LABEL:B4 ResNet18-NetVLAD performing slightly better when more candidates are recalled. LABEL:B2 RaPlace is worse than either of these at fewer recalled candidates but approaches LABEL:B3 Open-RadVLAD when more database candidates are retrieved.
VII-B Failure cases
There is in fact a precipitous drop in localisation success rate for all methods in LABEL:DA1-vs-LABEL:DA2 in twolochs. We attribute this to twolochs (Sec. III-D) being the longest route (there therefore being more candidate scans to match to – i.e. the search space being larger). In Fig. 5(b) this can be seen in the ground truth matrix (Top Right) by the large region to the right of the white strip, where the query trajectory ends at Loch Laggan and therefore only overlaps with the reference trajectory (Middle Right in Fig. 5(b)) for approximately of the route, with the reference trajectory continuing north-east all the way along both bodies of water in Loch Laggan. This may also be due to the long periods driven along either Loch Laggan or Lochan na h-Earba where there is a dearth of distinctive scenery in an entire hemisphere of the radar scan. An example scan from this type of scene can be seen in the third Cartesian frame in Fig. 5(b) (Bottom). Therefore, twolochs is a useful test setting for robust recognition with vast maps as well as with featureless scenery, as radar place recognition matures. For this challenging route, only radar-specific methods perform well and we have LABEL:B2 RaPlace performing best, with LABEL:B3 Open-RadVLAD matching it in Recall@10. LABEL:B4 ResNet18-NetVLAD performs poorly on this pair of trajectories, likely due to clustering of the reference trajectory features, many of which are irrelevant to the query trajectory.
VII-C Comparison against visual place recognition
Finally, Fig. 10 shows a comparison between visual place recognition and radar place recognition on this dataset. We see that radar has better performance in this task on this dataset across both experimental settings. For example, we have almost recall at any number of database candidates () – see the orange curve – while a vision-based approach has less than Recall@1 and this rises to only about (red curve). This was an experiment on LABEL:DB1 versus LABEL:DB2, with little appearance variation (see Fig. 7, top left) The next experiment (blue and green) is on LABEL:DM1 versus LABEL:DM2, which has a lot of appearance variation (see Fig. 7 where LABEL:DM2 is under very poor lighting – with driving taking place at night, in the dark). Here, radar does only very slightly worse – with over Recall@1 (Fig. 10 leftmost blue dot and Tab. III third row, fourth column). Performance with the camera, however, is drastically reduced – e.g. dropping to approximately Recall@1. This is an important result as it confirms that radar is somewhat immune to extreme appearance variation and is a useful counterpart (for localisation in this case) to the comparison to [32] where radar was shown to be quite robust to appearance change in urban environments in the task of vehicle detection and tracking.
VIII Conclusion
We have presented a novel radar dataset, carefully gathered under unique and challenging conditions. This dataset has been crafted to catalyse advancements in the emerging field of radar place recognition. We explored the use of our dataset in this task over a series of comprehensive experiments and evaluations, carried out across various open-source radar place recognition systems. This was not only to showcase the dataset but also to establish a robust platform for future research in this evolving area. We found that the best performing radar place recognition methods in terms of successful localisation queries over thousands of trials over pairs of sequences from distinct routes were the recently available open-source systems OpenRadVLAD, RaPlace, and ResNet18-NetVLAD the last of which is released alongside this dataset. We also proved the usefulness of the dataset in confirming appearance-change robustness for radar in comparison to visual sensors.
In the future, we would like to document new radar place recognition performance against this benchmark, and plan to do this online at e.g. paperswithcode.com/datasets.
Acknowledgements
[…]
We would also like to thank our partners at Navtech Radar, as well as Ian and Simon from Experience the Country and Stuart and Tony from Highland All Terrain, for expertly guiding and driving the vehicle around the data collection sites.